I'm really lost on this problem and there are no examples in my book. Please help!
z/(z+3) = 3z/(5z-1)

Answer by AnlytcPhil(1806)

I'm really lost on this problem and there are no examples in my book.
Please help!                

z/(z+3) = 3z/(5z-1)

[Notice that to type fractions all on one line, numerators and
 denominators must be placed within parentheses.  Or you don't
 have to use parentheses if you type on three lines like this:]
           

  z       3z
————— = ——————
 z+3     5z-1

The LCD is (z+3)(5z-1).  So multiply both sides by that

              z                   3z
(z+3)(5z-1)·————— = (z+3)(5z-1)·——————
             z+3                 5z-1

Simplify by canceling:

  1           z            1      3z
(z+3)(5z-1)·————— = (z+3)(5z-1)·——————
             z+3                 5z-1

          (5z-1)z = (z+3)3z

          z(5z-1) = 3z(z+3)

          5zē - z = 3zē + 9z

        2zē - 10z = 0

        2z(z - 5) = 0

Set each factor = 0.

Setting 2z = 0, gives z = 0

Setting z-5 = 0, gives z = 5      
         
Both check.

Edwin
AnlytcPhil@aol.com